OpenStudy (anonymous):

What is the equation (in slope-intercept form) of a line that passes through the point (3, -2) and has a slope of -2?

OpenStudy (anonymous):

y = mx + b, you know x and y from (3,-2) which is an (x, y) coordinate and you know m, which is slope, which is -2, now just find b

OpenStudy (anonymous):

plug, x y, and m into the equation y = mx + b and find b then plug b, m and x back into y = mx + b